Lina Chen invested $24,000, part at 8% and the rest at 7.2%. How much did she invest at each rate if her income from the 8% investment is two thirds of the 7.2% investment?
Answer by josmiceli(19441)

x = amount invested @ 8%
(24000 - x) = amount invested @ 7.2%
(.08)*x = 2/3(.072)(24000 - x)
3*(.08x) = 2*(.072)(24000 - x)
.24x = .144(24000 - x)
.24x = 3456 - .144x
.384x = 3456
x = 9000
(24000 - x) = 15000
check
(.08)*9000 = 2/3(.072)(24000 - 9000)
720 = 2/3(.072)(15000)
720 = (.144)(5000)
720 = 720